WUT IS 2 OVER N=5 OVER 10
Answer by checkley71(8403)

2/N=5/10 CROSS MULTIPLY
5N=2*10
5N=20 NOW DIVIDE BOTH SIDES BY 5.
X=20/5
X=4
PROOF:
2/4=5/10 CROSS MULTIPLY.
20=20
OR REDUCE.
1/2=1/2